OpenStudy (anonymous):

Given the point (5, 6) and the slope of 5, find y when x = 21.

OpenStudy (anonymous):

Apply the formula \[y -y_{1} = m(x-x_{1})\] where: \[m\]is the slope of the curve \[(x_{1}, y_{1})\] is your given point. Substitute those numbers in and make y the subject.

OpenStudy (anonymous):

It would be 6-y=21(x-5)

OpenStudy (anonymous):

@Azteck

OpenStudy (anonymous):

LHS is not quite correct. Please refer back to the formula I gave you compare the LHS of that formula to what you did on your LHS.

OpenStudy (anonymous):

Remember: \[y_{1}\]is the coordinate of your given point. Not y.

OpenStudy (anonymous):

y-6=21(5-x)?? I don't understand:(

OpenStudy (anonymous):

LHS means Left Hand side. sorry if I confused you with that. Your RHS (Right hand side) was correct before. Why did you change it?

OpenStudy (anonymous):

so I had to change just the y? like this? y-6=21(x-5)